# MGS-SLAM: Monocular Sparse Tracking and Gaussian Mapping with Depth Smooth Regularization

### v1

@@ III-C System Components
System initialization. Similar to DPVO, The system uses 8 frames for initialization. The pose of the new frame is initialized using a constant velocity motion model. Adding new patches and frames until 8 frames are accumulated and then running 12 iterations of the update operator. The 8 frames in the initialization are used as MVS network inputs to estimate the priori depth of the first frame. The backend uses the first priori depth as the foundation to initialize the Gaussian map.
Figure 5: Priori depth optimization. this optimization strategy in the SDAR is to correct the geometry of the priori depth map from MVS network and align the scale with the sparse point cloud map.
Keyframe selection. In the frontend tracking process, we always consider the 3 most recent frames as keyframes to fulfill the constant velocity motion model requirement. However, these 3 frames are not utilized for Gaussian mapping. Instead, we assess whether 4th frame satisfies Gaussian covisibility criteria. If it does, we add it into the mapping process in the backend; otherwise, we discard this frame. Between two keyframes $i$, $j$, we define the covisibility using Intersection of Union (IOU):
@@
where $\mathcal{G}_{i}$, $\mathcal{G}_{j}$ are visible Gaussians in the viewpoints of frame $i$ and frame $j$.
@@
Firstly, We use a sparse point cloud map with better geometric accuracy to correct the priori depth map from the MVS network estimate. Specifically, the sparse point cloud is transformed to the camera coordinate where the priori depth map is located, and then the sparse point cloud is projected to the sparse depth map. The priori depth map and the sparse depth map conform to the normal distribution of $\hat{D}_{d}\sim\mathcal{N}(\mu_{d},\sigma^{2}_{d})$ and $D_{s}\sim\mathcal{N}(\mu_{s},\sigma^{2}_{s})$. Align the priori depth map with the sparse depth map using the following equation:
@@
where $\hat{\sigma}_{d}$ and $\hat{\mu}_{d}$ are the mean and standard deviation status- tics of the sparsified priori depth map extracted from $\hat{D}_{d}$ at the pixel coordinates of $D_{s}$. This strategy corrects the prior depth errors and maintains the scale consistency, as shown in Fig. 5.
Secondly, we backproject the optimized prior depth map with RGB color into space, generating a new point cloud. Subsequently, downsampling is performed on this new point cloud. New Gaussians are then initialized with the downsampled point cloud as the center and added to the Gaussian map. This operation ensures consistent scale between the Gaussian map and the sparse point cloud map, while downsampling helps mitigate errors in the prior depth map.
@@
Finally, to achieve scale closure, we leverage the real-time rendering capability of the Gaussian map to generate the depth map of the frame being tracked at the frontend. We then initialize the depth of the tracking frame’s point cloud using this depth map. This strategy ensures that the frontend track aligns with the scale of the backend Gaussian map, facilitating the transfer of scale consistency from the backend to the frontend.
TABLE I: ATE [cm] results on Tum Dataset
Input Method fr1/desk fr2/xyz fr3/office Avg.
RGB-D iMAP 4.90 2.00 5.80 4.23
NICE-SLAM 4.26 6.19 6.87 5.77
Vox-Fusion 3.52 1.49 26.01 10.34
SplaTAM 3.35 1.24 5.16 3.25
Mono. DSO 22.40 1.10 9.50 11.00
DROID-VO 5.20 10.70 7.30 7.73
DPVO 3.80 0.54 7.00 3.78
MonoGS 4.15 4.79 4.39 4.44
Ours 2.33 0.44 3.00 1.92
@@
RGB-D iMAP 3.12 2.54 2.31 1.69 1.03 3.99 4.05 1.93 2.58
NICE-SLAM 0.97 1.31 1.07 0.88 1.00 1.06 1.10 1.13 1.07
Vox-Fusion 1.37 4.70 1.47 8.48 2.04 2.58 1.11 2.94 3.09
@@
IV EXPERIMENTS
We evaluate our proposed system on a series of real and synthetic datasets, including the TUM dataset [29] and the Replica dataset [24]. We compare the pose estimation accuracy (ATE) and Novel View Rendering quality with previous works, utilizing experimental results from papers or open-source code of these works. The experimental data from the source code represents the average of three runs. Additionally, we conduct an ablation study to demonstrate the effectiveness of the components of our system. Finally, we analyze the system runtime.
@@ IV-A Implementation Details
We evaluate our proposed system and other methods on a desktop with an Intel Core i7 12700 processor running at 3.60GHz and a single NVIDIA GeForce RTX 3090. The size of input images is consistent with the dataset size in our system, but these images are resized to 512$\times$384 when estimating prior depth maps to better align with the resolution of MVS training images and enhance efficiency. Similar to Gaussian Splatting, mapping rasterization and gradient computations are implemented using CUDA. The remainder of our system pipeline is developed with PyTorch. For map optimization, we set the maximum gradient threshold to 0.0002 and the minimum opacity threshold to 0.65 for the Gaussians in the densify and prune operation of the mapping.
@@ IV-B Camera Tracking Accuracy
For camera tracking accuracy, we report the Root Mean Square Error (RMSE) of the keyframes Absolute Trajectory Error (ATE). We benchmark our system against other approaches that, like ours, do not have explicit loop closure. We compared not only the monocular methods but also the RGB-D methods. The comparative work is very comprehensive including traditional direct visual odometry [9], learning-based SLAM[10], neural implicit slam [26], [30], and more recently Gaussian Splatting-based SLAM [22], [23].
Tab. I shows the tracking results on the TUM dataset. The tracking accuracy of SplaTAM exceeds other RGB-D methods. The tracking accuracy of our system also outperforms other monocular methods by 40%, while surpassing SplaTAM. Tab. II shows the tracking results on the Replica dataset. The same Gaussian Splatting-based MonoGS with monocular input experienced tracking loss in some scenes in the Replica dataset. Our system completes the full scene tracking, demonstrates robustness, and achieves state-of-the-art tracking results. In addition, The experimental data from both tables show that our tracking performance is superior to the DPVO on which the frontend is based. This demonstrates the effectiveness of our combination of sparse visual odometry and Gaussian mapping to achieve a more robust and accurate SLAM system.
TABLE III: Rendering performance on Replica Dataset. Best results are highlighted as first, second, and third
Method Metric R0 R1 R2 O0 O1 O2 O3 O4 Avg.
NICE- SLAM PSNR[dB]$\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
[…]
GO- SLAM PSNR[dB]$\uparrow$ 23.25 20.70 21.08 21.44 22.59 22.33 22.19 22.76 22.04
@@
TABLE IV: Mapping losses ablation on Office 0
@@
TABLE V: Sparse-Dense Adjustment Ring ablation on Office 0
@@ IV-C Novel View Rendering
We evaluated the methods for novel view rendering on Replica. To evaluate map quality, we report standard photometric rendering quality metrics (PSNR, SSIM and LPIPS) following the evaluation protocol used in [31]. The methods we are comparing have RGB-D input and monocular input. NICE-SLAM and Vox-Fusion are RGB-D input and the rest are monocular input. The rendering results of MonoGS is the average before it tracking loss. Tab. IV-B shows the results, our proposed system performs state-of-the-art in most scenes. The visualisation of the rendering is shown in Fig. 6, where the quality of our rendered image is higher than the other methods and almost indistinguishable from the ground truth.
IV-D Ablative Analysis
Mapping losses ablation. We changed the losses function of the original 3D Gaussian Splatting by introducing depth loss, smooth loss, and isotropic loss. As shown in Tab. IV, we did an ablation study of these losses and evaluated their impact through ATE, PSNR, and Depth L1 metrics. The results show that all these losses contribute to the accuracy improvement of the system. It is worth noting that the depth loss utilizes the estimated priori depth map from the Network. Although this improves the geometry of our system, it also leads to incorrect geometric guidance. After adding the smooth loss, the geometric accuracy is further enhanced, proving that the smooth loss can correct some adverse effects of the estimated priori depth map.
TABLE VI: Runtime analysis on TUM fr3/office
Method Tracking/Frame[s] Mapping/Frame[s] Total Time[s] FPS
SplaTAM 3.65 1.44 10652 0.24
MonoGS 0.55 1.70$\sim$2.00 798.90 3.20
Ours 0.25 1.61$\sim$1.90 628.1 4.00
@@ IV-D Ablative Analysis
Sparse-Dense Adjustment Ring ablation. We propose the Sparse-Dense Adjustment Ring (SDAR) strategy to unify the frontend and backend scales. This strategy comprises three components (Sec. III-C). We conducted an ablation study of these three components to demonstrate their effect on the system. The metrics for the ablation study are also adopted as PSNR, SSIM, and LPIPS. As shown in Tab. V, the contribution of SDAR to the system is mainly in the tracking accuracy ATE. The tracking accuracy of the system is similar to DPVO without the SDAR strategy. The tracking accuracy of the system is state-of-the-art with the SDAR strategy. This demonstrates the significance of combining sparse visual odometry and Gaussian mapping to build a more accurate and robust SLAM system.
MVS window analysis. Since the MVS network requires at least one reference frame and one source frame, the minimum window size is 2, and the maximum is the number of keyframes required for DPVO initialization. As depicted in Fig. 7, we have analyzed the effect of different window sizes of MVS on the accuracy and speed of the system. Since our MVS network consists of 2D convolutions, increasing the window size has little effect on inference time. However, increasing the window size improves the system’s tracking accuracy and mapping quality. This is because more keyframes with different views provide additional geometrical cues, which can estimate better prior depth map and enhance system accuracy.
IV-E Runtime Analysis
As shown in Tab. VI, we analyzed the runtime of our proposed system on the TUM fr3/office dataset and compared it to other methods. The results indicate we achieve higher frames per second (fps) than other Gaussian Splatting-based SLAM methods. This advantage is primarily attributed to the fast frontend tracking. These findings underscore the effectiveness of our approach, which combines sparse visual odometry and Gaussian mapping.
@@ V CONCLUSIONS
This letter introduces MGS-SLAM, a novel Gaussian Splatting-based SLAM framework. For the first time, our framework integrates sparse visual odometry with Gaussian mapping, addressing the tracking loss issue in larger scenes encountered by Gaussian Splatting-based SLAM and eliminating the need for depth maps input. We develop a lightweight MVS depth estimation network to facilitate this integration. Additionally, we propose the Sparse-Dense Adjustment Ring (SDAR) strategy to adjust the scale between the sparse map and the Gaussian map. Comparative evaluations demonstrate that our approach achieves state-of-the-art accuracy compared to previous methods.

### v2

@@ III-C System Components
System initialization. Similar to DPVO, The system uses 8 frames for initialization. The pose of the new frame is initialized using a constant velocity motion model. We add new patches and frames until 8 frames have been accumulated, and then run 12 iterations of the update operator. The 8 frames in the initialization are used as MVS network inputs to estimate the priori depth of the first frame. The backend uses the first priori depth as the foundation to initialize the Gaussian map.
Figure 5: Priori depth optimization. this optimization strategy in the SDAR is to correct the geometry of the priori depth map from the MVS network and align the scale with the sparse point cloud map.
Keyframe selection. In the frontend tracking process, we always consider the 3 most recent frames as keyframes to fulfill the constant velocity motion model requirement. However, these 3 frames are not utilized for Gaussian mapping. Instead, we assess whether 4th frame satisfies Gaussian co-visibility criteria. If it does, we add it to the mapping process in the backend; otherwise, we discard this frame. This method can determine whether the tracked frame has new information exceeding a threshold, improve the efficiency of keyframe usage, and reduce memory consumption. Between two keyframes $i$, $j$, we define the co-visibility using Intersection of Union (IOU):
@@
where $\mathcal{G}_{i}$, $\mathcal{G}_{j}$ are visible Gaussians in the viewpoints of frame $i$ and frame $j$. If IOU is less than a threshold, the system will create a new keyframe.
@@
Firstly, We use a sparse point cloud map with better geometric accuracy to correct the priori depth map from the MVS network estimate. The priori depth map and the sparse depth map conform to the normal distribution of $\hat{D}_{d}\sim\mathcal{N}(\mu_{d},\sigma^{2}_{d})$ and $D_{s}\sim\mathcal{N}(\mu_{s},\sigma^{2}_{s})$. Align the priori depth map with the sparse depth map using the following equation:
@@
where $\hat{\mu}_{d}$ and $\hat{\sigma}_{d}$ are the mean and standard deviation statustics of the sparsified priori depth map extracted from $\hat{D}_{d}$ at the pixel coordinates of $D_{s}$. This strategy corrects the prior depth errors, as shown in Fig. 5.
@@
Secondly, we backproject the optimized prior depth map with RGB color into space, generating a new point cloud. Subsequently, downsampling is performed on this new point cloud. New Gaussians are then initialized with the downsampled point cloud and added to the Gaussian map.
Finally, to achieve scale closure, we leverage the real-time rendering capability of the Gaussian map to generate the depth map of the frame being tracked at the frontend. We then initialize the depth of the tracking frame’s point cloud using this depth map. This strategy ensures that the frontend track aligns with the scale of the backend Gaussian map.
IV Experiments
We evaluate our proposed system on a series of real and synthetic datasets, including the TUM dataset [26], Replica dataset [27] and ICL-NUIM dataset [28]. We compare the pose estimation accuracy (ATE), novel view rendering quality and geometric reconstruction quality with previous works, utilizing experimental results from papers or open-source code of these works. The experimental data from the source code represents the average of three runs. Additionally, we conduct some ablation studies to demonstrate the effectiveness of our system’s components. Finally, we analyze the system runtime and memory.
TABLE I: ATE [cm] results on TUM Dataset.
Input Method fr1/desk fr1/desk2 fr1/plant fr2/xyz fr3/office Avg.
RGB-D SplaTAM 3.35 6.54 2.74 1.24 5.16 3.81
Co-SLAM 2.70 4.31 4.74 1.90 2.60 3.25
ESLAM 2.30 3.78 2.11 1.10 2.40 2.34
Mono. DSO 22.40 91.60 12.10 1.10 9.50 27.34
DROID-VO 5.20 9.90 2.80 10.70 7.30 7.18
MonoGS 4.15 7.16 7.82 4.79 4.39 5.66
Photo-SLAM 1.54 21.00 3.67 0.98 1.26 5.69
DPVO 3.80 6.40 4.70 0.54 7.00 4.49
Ours 2.33 5.32 3.55 0.44 3.00 2.93
@@ IV Experiments
RGB-D Co-SLAM 0.70 0.95 1.35 0.59 0.55 2.03 1.56 0.72 1.06
ESLAM 0.71 0.70 0.52 0.57 0.55 0.58 0.72 0.63 0.62
NeSLAM 0.60 0.93 0.52 0.41 0.43 0.57 0.96 0.83 0.66
GS-SLAM 0.48 0.53 0.33 0.52 0.41 0.59 0.46 0.70 0.50
@@
TABLE III: ATE [cm] results on ICL-NUIM Dataset
Input Method L0 L1 L2 L3 O0 O1 O2 O3 Avg.
RGB-D Co-SLAM 1.15 0.85 1.03 16.46 52.46 3.60 1.76 39.15 14.56
ESLAM 0.45 0.49 1.61 5.84 0.42 1.37 1.01 0.46 1.46
SplaTAM 0.53 0.70 1.13 4.63 0.42 1.03 0.92 1.16 1.32
Mono. DSO 1.00 2.00 6.00 3.00 21.00 83.00 36.00 64.00 27.00
DROID-VO 1.00 12.30 7.20 3.20 9.50 4.10 84.20 50.40 21.49
MonoGS 6.40 21.21 31.40 100.76 13.87 35.76 24.73 73.42 38.44
Photo-SLAM 0.54 4.52 0.72 0.98 3.41 18.19 1.54 4.71 4.33
DPVO 0.60 0.60 2.30 1.00 6.70 1.20 1.70 63.50 9.70
Ours 0.58 0.50 1.82 0.77 1.46 1.01 1.19 1.49 1.10
@@ IV-A Implementation Details
We evaluate our proposed system and other methods on a desktop with an Intel Core i7 12700 processor running at 3.60GHz and a single NVIDIA GeForce RTX 3090. The size of input images is consistent with the dataset size in our system. Similar to Gaussian Splatting, mapping rasterization and gradient computations are implemented using CUDA. The remainder of our system pipeline is developed with PyTorch. For map optimization, we set the maximum gradient threshold to 0.0002 and the minimum opacity threshold to 0.65 for the Gaussians in the densify and prune operation.
@@ IV-B Camera Tracking Accuracy
For camera tracking accuracy, we report the Root Mean Square Error (RMSE) of the keyframe’s Absolute Trajectory Error (ATE). We benchmark our system against other approaches. The comparative works are very comprehensive including traditional visual odometry DSO [29], learning-based visual odometry DROID-VO [7], neural implicit-based NICER-SLAM [30], NeSLAM [25], ESLAM [31], Co-SLAM [32] and more recently Gaussian Splatting-based SplaTAM [16], MonoGS [20], GS-SLAM [17], Photo-SLAM [21].
Tab. I shows the tracking results on the TUM dataset. The tracking accuracy of our system outperforms other monocular methods by 35% and is comparable to ESLAM using RGB-D input. Tab. II and Tab. III show that our system achieved the best tracking performance compared to other systems including monocular and RGB-D. In addition, The experimental data from the tables show that our tracking performance is superior to the DPVO on which the frontend is based. This demonstrates the effectiveness of our combination of sparse visual odometry and Gaussian mapping in achieving a more robust and accurate SLAM system.
TABLE IV: Rendering performance on Replica Dataset. Best results are highlighted as first, second, and third
Input Method Metric R0 R1 R2 O0 O1 O2 O3 O4 Avg.
RGB-D NICE- SLAM PSNR[dB]$\uparrow$ 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 24.42
@@
Mono. GO- SLAM PSNR[dB]$\uparrow$ 23.25 20.70 21.08 21.44 22.59 22.33 22.19 22.76 22.04
[…]
TABLE V: Reconstruction performance on Replica Dataset. Best results are highlighted as first, second, and third
Input Method Depth L1[cm]$\downarrow$ Acc.[cm]$\downarrow$ Comp.[cm]$\downarrow$ Comp. Ratio[$<$5cm]$\uparrow$
Mono. MonoGS 36.58 74.02 19.30 37.51
Photo-SLAM 19.73 53.70 8.08 49.46
GO-SLAM 4.39 3.81 4.79 78.00
NICER-SLAM - 3.65 4.16 79.37
Ours 7.77 7.51 3.64 82.71
TABLE VI: Mapping losses ablation on Office 0
@@
TABLE VII: Sparse-Dense Adjustment Ring ablation on Office 0
@@ IV-C Novel View Rendering
We evaluated the methods for novel view rendering on Replica. To evaluate map quality, we report standard photometric rendering quality metrics (PSNR, SSIM and LPIPS). The methods we are comparing have RGB-D input and monocular input. NICE-SLAM [12], Vox-Fusion [33], ESLAM [31] and Co-SLAM [32] are neural implicit-based RGB-D input and the rest are monocular input. We take the average of frames other than keyframes to evaluate rendering quality. Tab. IV-B shows the results, our proposed system performs state-of-the-art in most scenes. The visualization of the rendering is shown in Fig. 6, where the quality of our rendered image is higher than the other methods and almost indistinguishable from the ground truth.
IV-D Geometric Reconstruction
We evaluated the methods for geometric reconstruction on Replica. The methods evaluated are all monocular differentiable rendering SLAM approaches. We report standard mesh geometric reconstruction metrics (Depth L1, Accuracy, Completion, Completion Ratio). Tab. V shows the results, our method achieved the best results in terms of Completion and Completion Ratio metrics. It is worth noting that our geometric reconstruction performance is 50% higher than other monocular 3D Gaussian Splatting-based SLAM, which proves the effectiveness of our method in utilizing the MVS network to promote geometric reconstruction. Furthermore, this better geometric reconstruction also improves the rendering.
IV-E Ablative Analysis
Mapping losses ablation. We changed the loss function of the vanilla 3D Gaussian Splatting by introducing depth loss, smooth loss, and isotropic loss. As shown in Tab. VI, we did an ablation study of these losses. The results show that all these losses contribute to the accuracy improvement of the system. It is worth noting that the incorrect geometric guidance caused by the depth loss using the prior depth maps was corrected after adding depth smoothing loss.
TABLE VIII: Runtime and Memory analysis on TUM and Replica Datasets
[…]
TUM SplaTAM 14.28ms 16.77ms 2.85s 0.50s 526.32 42.31MB
MonoGS 6.78ms 12.67ms 0.65s 1.90s 1126.10 2.80MB
Photo-SLAM - 8.91ms 33.33ms - 1648.20 12.77MB
Ours - 11.90ms 35.17ms 1.85s 1173.21 1.96MB
Replica SplaTAM 25.43ms 23.80ms 2.25s 1.43s 125.64 273.09MB
MonoGS 10.78ms 20.50ms 1.10s 3.07s 769.00 24.50MB
Photo-SLAM - 15.18ms 37.45ms - 911.26 22.21MB
Ours - 18.98ms 38.41ms 2.97s 776.50 20.90MB
@@ IV-E Ablative Analysis
Sparse-Dense Adjustment Ring ablation. We propose the Sparse-Dense Adjustment Ring (SDAR) strategy to unify the frontend and backend scales. This strategy comprises three components (Sec. III-C). We conducted an ablation study of these three components to demonstrate their effect on the system. As shown in Tab. VII, the contribution of SDAR to the system is mainly in the tracking accuracy ATE. The tracking accuracy of the system is similar to DPVO without the SDAR strategy.
MVS window analysis. As depicted in Fig. 7, we have analyzed the effect of different window sizes of MVS on the accuracy and speed of the system. Since our MVS network consists of 2D convolutions, increasing the window size has little effect on inference time. However, increasing the window size improves the system’s tracking accuracy and mapping quality. This is because more keyframes with different views provide additional geometrical cues.
IV-F Runtime and Memory Analysis
As shown in Tab. VIII, We quoted the method [18] to analyze the runtime and memory of our system and compare it to other methods on the TUM and Replica datasets. The memory is the memory usage of the checkpoint. Some methods do not use this metric and are represented by shorter lines. The metric of tracking each frame contains the inference time of the MVS network in our method, and the MVS network runs on keyframes. The results show that our tracking speed is similar to Photo-SLAM. However, our method achieved better geometry at the expense of tracking time, resulting in more compact checkpoint and the best memory utilization.
@@ V Conclusions
This letter introduces MGS-SLAM, a novel Gaussian Splatting-based SLAM framework. For the first time, our framework jointly optimizes sparse visual odometry tracking and 3D Gaussian mapping, enhancing tracking accuracy and geometric reconstruction precision of Gaussian maps when only given RGB image input. We develop a lightweight MVS depth estimation network to facilitate this integration. Additionally, we propose the Sparse-Dense Adjustment Ring (SDAR) strategy to adjust the scale between the sparse map and the Gaussian map. Comparative evaluations demonstrate that our approach achieves state-of-the-art accuracy compared to previous methods. We believe that this innovative method will bring some inspiration to future works.